Hannah made a bank loan worth Php 200,550 with 5% interest compounded semi-annually (2 times in a year). If she is to pay at the end of the period only, how much will be the value of the loan after 2 years?


1
Expert's answer
2022-01-31T12:07:45-0500

A= P ( 1 + i/n)nt

A = Future Value

P = Principal Amount

i = interest rate

n = Number of times compunded in a year

t = time in years

A= P ( 1 + i/n)nt

P = 200,550

i = 5% p.a

n = 2 times

t = 2 years

A = 200,550 ( 1 + 0.05/2)2*2

A = 200,550 ( 1 + 0.025)4

A = 221,370

Value of Loan After 2 years = Php 221,370